OpenStudy (anonymous):

How do I find the integral of x^5(sqrt(4-x^3))

OpenStudy (luigi0210):

This? \(\Large \int x^5\sqrt{4-x^3}~dx\)?

OpenStudy (anonymous):

Yes! thanks

OpenStudy (luigi0210):

What have you tried so far?

OpenStudy (anonymous):

\[du = \frac{ 3x }{ 2\sqrt{4-x ^{3}} } dx\] \[du =-3x ^{2}dx\] \[du =5x ^{4}dx\]

OpenStudy (luigi0210):

@amistre64 could help you :3

OpenStudy (amistre64):

gonna have to do u sub and by parts together

OpenStudy (amistre64):

let u = 4-x^3 du = -3x^2 dx -du/3 = x^2 dx \[ \int x^3\sqrt{u}~du\] well, x^3 = 4-u \[ \int (4-u)\sqrt{u}~du\] by parts wont be needed, just a few powers is all

OpenStudy (amistre64):

forgot the -1/3 .... but thats the basic process id do

OpenStudy (anonymous):

How did you turn x^5 to 4-u when x^3=4-u

OpenStudy (phi):

amistre let \[ u = 4-x^3 \\ du = -3x^2 dx \] so \[ dx = -\frac{1}{3} \frac{du}{x^2} \] and \[ \int x^5 \sqrt{4-x^3} \ dx\rightarrow \int x^5 \sqrt{u}\left(-\frac{1}{3}\right) \frac{du}{x^2} \] which simplifies to \[ -\frac{1}{3} \int \frac{x^5}{x^2} \sqrt{u} \ du \\ -\frac{1}{3} \int x^3 \sqrt{u} \ du \] using \[ u = 4-x^3 \] and solving for x^3 we get \[ x^3 = 4-u \] and fnally \[ -\frac{1}{3} \int (4-u) \sqrt{u} \ du \\ -\frac{1}{3} \int 4u^\frac{1}{2}-u^\frac{3}{2} \ du \]

OpenStudy (anonymous):

Oh wow ofcourse. Thanks a lot all. I really appreciate this effort!

OpenStudy (amistre64):

:)